OpenStudy (anonymous):

what is 4^(3/2)? How do you show work for it?

OpenStudy (luigi0210):

It's the same as: \[\sqrt[2]{4^3}\]

OpenStudy (luigi0210):

What's 4^3?

OpenStudy (anonymous):

oh alright, 4^3=64...

OpenStudy (klimenkov):

What is \(\sqrt 4\) ?

OpenStudy (luigi0210):

Yes, now what's \[\sqrt{64}\]

OpenStudy (luigi0210):

@klimenkov 's method would of worked too

OpenStudy (anonymous):

8

OpenStudy (luigi0210):

And that's your answer

OpenStudy (anonymous):

alrighty what is 8^(2/3)?

OpenStudy (luigi0210):

what's 8^2?

OpenStudy (anonymous):

64

OpenStudy (luigi0210):

and what's \[\sqrt[3]{64}\]

OpenStudy (anonymous):

ohhhh 4

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

thanks dood

OpenStudy (klimenkov):

I think i would be easier to compute \(8^{\frac13}\) first.

OpenStudy (anonymous):

ohh dood 1 more please

OpenStudy (anonymous):

what is √(8)/2^(1/2)?

OpenStudy (luigi0210):

Is that 1/2 with the whole equation or just 2?

OpenStudy (anonymous):

sq root 8 divided by 2 to the half power

OpenStudy (luigi0210):

Well the 1/2 makes that 2 a rad 2: \[\frac{\sqrt{8}}{\sqrt{2}}\] Now what does \[\sqrt{8}\] simplify into?

OpenStudy (anonymous):

2√(2)

OpenStudy (luigi0210):

Yup! \[\frac{2\sqrt{2}}{\sqrt{2}}\] Can you finish it?

OpenStudy (anonymous):

2, right?

OpenStudy (luigi0210):

Mhm :)

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

do you know how to calculate average rate of change?

OpenStudy (luigi0210):

Anytime. Average rate of change? As in calculus?

OpenStudy (anonymous):

yes :D

OpenStudy (luigi0210):

Sorry, I haven't had calculus for a while :/ just open a new question and call Nincompoop, he's great at calculus

OpenStudy (anonymous):

ah alrighty thanks dood.

OpenStudy (luigi0210):

Yup, and if he's not on call dan815

OpenStudy (anonymous):

kk